5 years ago I was thrice as old as my son and 10 years later I shall be twice as old as my son. How old am I now?
Answer by jorel1380(3719)

Me-5=3(Son-5)
Me+10=2(Son+10)
--------------
Me-5=3Son-15
Me+10=2Son+20
Me=3Son-10
3Son-10+10=2Son+20
3Son=2Son+20
Son=20
Me=50.